If given:
A=+68+ a = ?
B = ? b = +8+
C = ? c =+5+
What is the value of a?
A.) +7.68+
B.) +9.31+
C.) +11.22+
D.) +13.18+

Answer by Boreal(15235)

Law of Cosines
a^2=b^2+c^2-2bc cos A
=64+25-2(40) *0.3746=89-29.97=59.03
sqrt(59.03)=7.68
A.